# Irreducible finite-index representations of automorphism groups of trees

### v1

@@ 3 Restricting to spherical representations
for all $g\in G,\ v,w\in H$. This means that $A$ commutes with the representation, i. e. $A\in\mathrm{End}(\pi)$. By the Schur Lemma ([2], Lemma 3.21), $\mathrm{End}(\pi)$ is a division algebra over $\mathbb{R}$. If we fix a nonzero $v\in H^{K}$ the map $\mathrm{End}(\pi)\rightarrow H^{K}$ defined by $\phi\mapsto\phi(v)$ is injective, since $\pi$ is irreducible, thus $\dim(\mathrm{End}(\pi))\leq\dim(H^{K})$. Since $\dim(H^{K})$ is finite, so is $\dim(\mathrm{End}(\pi))$, and it is well known that then $\mathrm{End}(\pi)$ must be isomorphic to $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$.
If $\mathrm{End}(\pi)$ is isomorphic to $\mathbb{R}$, then we are done. In the other two cases, we have $A=\alpha Id+\beta A_{1}$ with $\alpha,\beta\in\mathbb{R}$ and $A_{1}^{2}=-Id$. We wish to prove that $\beta=0$. Since $A$ and $\alpha Id$ are self-adjoint and both belong to $\mathrm{End}(\pi)$, the same must be true for $\beta A_{1}$ and therefore, if $\beta\neq 0$, also for $A_{1}$. Then for every $v\in H$ we have
[…]
4 An explicit form for spherical representations
The space $C(S)$ is the space of $K$-right-invariant functions on $G$ that are left-invariant under some open compact subgroup. We can identify this space with the space of functions on $G/K$ (the set of right cosets of $G$ under $K$) that are left-invariant under some open compact subgroup of $G$. The set $G/K$ can be identified with the orbit of $o$ in $T$ under the action of $G$, which is either the full vertex set $V$ or the set of vertices at even distance from $o$ (Proposition 2.2(c)).
In this section, we assume that we have the full vertex set $V$, i. e. that the action of $G$ is transitive on the vertices. In this case, the tree must be homogeneous. In the last section of the paper, we will extend the results to the other case.
Proposition 3.2 now tells us that we can see the representation $(\pi,H)$ as a subrepresentation of $(\Pi,C)$, where $C$ is the space of functions on $V$ that are left-invariant under some open compact subgroup of $G$ and $\Pi$ is the left action of $G$ on $C$. Therefore, from now on we assume that $H\subset C$ and $\pi$ is the restriction of $\Pi$ to $H$.
It is known that $(G,K)$ is a Gel’fand pair, which means that the convolution algebra of compactly supported complex functions on $K\backslash G/K$ is commutative. This follows ([1], Proposition 7) from the stronger fact that for every $g\in G$ we have $g^{-1}\in KgK$: the set $KgK$ is the set of $h\in G$ such that $d(ho,o)=d(go,o)$, and we have $d(g^{-1}o,o)=d(o,go)$.
Proposition 4.1.
For any irreducible representation $(\pi,H)$ of $G$ that preserves a sesquilinear form of finite index $q$, the space $H^{K}$ of $K$-fixed vectors is at most 111-dimensional.
Proof.
We consider the convolution algebra $\mathcal{A}$ of compactly supported functions on $K\backslash G/K$ (i. e. compactly supported functions on $G$ which are invariant under $K$ on the left and on the right). We consider real functions if $\mathbb{K}=\mathbb{R}$, and complex functions if $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{H}$.
@@ 4 An explicit form for spherical representations
To prove our proposition, we follow the proof of Proposition 5.4 in [11]. As in that proof, we observe that this algebra is commutative and symmetric, the latter meaning that the adjoint of every element of $\mathcal{A}$ (with respect to the sesquilinear form preserved by $\pi$) is again in $\mathcal{A}$. Since in our case $g^{-1}\in KgK$ for every $g\in G$, a stronger statement holds: every element of $\mathcal{A}$ is self-adjoint.
@@
As in [11] we see that $\mathcal{A}$ acts irreducibly on $H^{K}$ (which we assume to be non-trivial), we see that the restriction of the sesquilinear form on $H$ to $H^{K}$ is non-degenerate (say of index $q^{\prime}\geq 0$), and we use the fact, due to Naimark ([13], Corollary 2) that a commutative symmetric algebra of operators of a complex Hilbert space endowed with a sesquilinear form of finite index $q^{\prime}\geq 1$ preserves a non-negative subspace $L$ of finite dimension. This is also true in the case $q^{\prime}=0$, due to the Schur Lemma, as explained in the addendum to [11].
@@
In the case $\mathbb{K}=\mathbb{C}$, the result in [13] applied to the representation of $\mathcal{A}$ on $H^{K}$ ensures that $\mathcal{A}$ preserves a finite-dimensional subspace $L$ of $H^{K}$, and as the representation is irreducible it must be $L=H^{K}$. Since $\mathcal{A}$ is commutative every element of $\mathcal{A}$ is an endomorphism of the representation, so the Schur Lemma implies that $\mathcal{A}\cong\mathbb{C}$ and therefore $H^{K}$ is 111-dimensional.
In the case $\mathbb{K}=\mathbb{H}$, we apply the result in [13] to the complex representation of $\mathcal{A}$ on $H^{K}$ seen as a complex vector space and find that it preserves a finite-dimensional complex subspace $L$. The quaternionic vector space that it generates is finite-dimensional and must coincide with $H^{K}$ by irreducibility. By applying the Schur Lemma, we see again that $\mathcal{A}\cong\mathbb{C}$ and therefore $H^{K}$ must be 111-dimensional.
@@
In the case $\mathbb{K}=\mathbb{R}$, we apply the result in [13] to the complexification of the representation of $\mathcal{A}$ on $H^{K}$. We find that there exists a finite-dimensional complex vector space $L$ preserved under $\mathcal{A}\otimes\mathbb{C}$ and thus under $\mathcal{A}$. This implies that there exists a finite-dimensional real vector space $L^{\prime}\subseteq H^{K}$ preserved under $\mathcal{A}$, which must coincide with $H^{K}$ by irreducibility. The representation preserves a positive definite bilinear form on $L^{\prime}=H^{K}$, since the restriction of the sesquilinear form to $L$ is non-negative and non-degenerate. If we apply the Schur Lemma, we see that $\mathcal{A}$ is a finite-dimensional commutative division algebra over $\mathbb{R}$ and therefore it is isomorphic to $\mathbb{R}$ or $\mathbb{C}$. In the latter case, there exists $A\in\mathcal{A}$ such that $A^{2}=-Id$, which implies, since each $A\in\mathcal{A}$ is self-adjoint, that for every $v\in H^{K}$ it holds
@@
contradicting the fact that the form is positive definite. Therefore, it must be $\mathcal{A}\cong\mathbb{R}$, which implies that $H^{K}$ is 111-dimensional.
Proposition 4.1 implies that there exists a function $f\in H$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$.
@@
The function $f$ is an eigenfunction of the Laplacian, i. e. there exists some $\alpha\in\mathbb{K}$ such that $Lf=\alpha f$.
@@
Since the space of radial functions in $H$ (which is the same as the space of $K$-invariant functions in $H$) is 111-dimensional, and both $f$ and $Af$ are in $H$, we must have $Af=\alpha f$, and hence $Lf=\alpha f$, for some $\alpha\in\mathbb{K}$.
It is immediate to see, from Observation 4.4, that $f$ is uniquely determined (up to a scalar multiple) by the property $Lf=\alpha f$ and the fact that it is radial. We also observe that $f(0)\neq 0$, because otherwise $f$ would be identically zero.
[…]
We are left with the general case $\alpha\neq\pm 1$. From Observation 4.4 and $Lf=\alpha f$ we get $f(1)=\alpha f(0)$ and $f(2)=\frac{1}{d-1}(d\alpha f(1)-f(0))=\frac{(\alpha^{2}d)-1}{d-1}f(0)$. This implies $f(2)\neq f(0)$: otherwise, we would have $(\alpha^{2}d)-1=d-1$ and thus $\alpha=\pm 1$.
@@
For any irreducible spherical representation $(\pi,H)$ of $G$ that is not 111-dimensional, there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that the representation is equivalent to the left action of $G$ on the subspace of $C$ defined by $\{h\in C:Lh=\alpha h\}$. Conversely, for every $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, there is an irreducible spherical representation of $G$ that results from the construction.
@@
We start by picking a general irreducible spherical representation of $G$ and using the results above. We know that there exist $\alpha\in\mathbb{K}$ and $f\in H\subset C$ such that $Lf=\alpha f$ and the subspace of $K$-invariant functions in $H$ is the space of scalar multiples of $f$. We can assume $\alpha\neq\pm 1$, and thus $f(0)\neq f(2)$: otherwise, as we have seen, the representation is 111-dimensional.
Since the representation is irreducible, $H$ is generated (as a vector space) by $\{\pi(g)f:g\in G\}$. A function $\pi(g)f$ is radial with respect to $go$ and satisfies $L\pi(g)f=\alpha\pi(g)f$. The latter property is preserved under linear combinations, so we have $H\subseteq\{h\in C:Lh=\alpha h\}$.
To obtain the other inclusion, we pick a generic $h\in C$ such that $Lh=\alpha h$ and we show that it is in $H$. Since $H$ is $G$-invariant and $f\in H$, it is enough to show that we can express $h$ as a finite linear combination of translates (by elements of $G$) of $f$.
For every $v\in V$, let $N(v)$ be the set of neighbors of $v$ which are further away from the origin than $v$ (if $v=o$ we take the full set of neighbors). To build $h$ as a finite linear combination of translates of $f$, we start with the function $F=h(o)f(0)^{-1}f$, that agrees with $h$ at the origin. As $LF(o)=\alpha F(o)=\alpha h(o)=Lh(o)$, the sum of $F$ over $N(o)$ agrees with $h$. To make the two functions agree at every vertex in $N(o)$, we add to $F$ the function
$\sum_{w\in N(o)}{(h(w)-F(w))(f(0)-f(2))^{-1}\pi(g_{w})f},$
where $g_{w}\in G$ sends $o$ to $w$. The sum of the coefficients is zero, and this guarantees that $F$ is unchanged at $o$. At every $w\in N(o)$ the function changes by
$(h(w)-F(w))(f(0)-f(2))^{-1}f(0)$
@@
$\sum_{z\in N(o)\setminus\{w\}}{(h(z)-F(z))(f(0)-f(2))^{-1}f(2)}=-(h(w)-F(w))(f(0)-f(2))^{-1}f(2)$
due to the contribution from $\pi(g_{z})f$ for $z\neq w$. This adds up to $h(w)-F(w)$, so we have a new function $F$ that agrees with $H$ on the ball $B_{1}(o)$.
We can move forward by applying the same procedure to each of the sets $N(w)$ with $w\in N(o)$: this will allow to update the function $F$ so that it agrees with $h$ inside $N(w)$ without changing the values inside the ball $B_{1}(o)$, and finally get a function that agrees with $h$ on the ball $B_{2}(o)$. We can continue inductively with $B_{3}(o)$ and so on. After finitely many steps $F$ and $h$ will agree everywhere: $h$ is left-invariant under some open compact subgroup that contains $G(B_{n}(o))$ for some $n$ (Proposition 2.2(a)), so it is uniquely determined by the property $Lh=\alpha h$ and the values inside $B_{n}(o)$, which means that it suffices to have $F$ and $h$ agree on $B_{n}(o)$.
We have proved that the space $H$ coincides with $\{h\in C:Lh=\alpha h\}$, so the representation we started from is equivalent to the left action of $G$ on that space.
For the converse, we wish to prove that if we start from a generic $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, and consider the space $\{h\in C:Lh=\alpha h\}$ with the left action of $G$, we have a spherical irreducible representation. We know that there is a radial (hence $K$-invariant) nonzero function in this space, so we are only left to prove irreducibility. Suppose there is a nonzero $G$-invariant subspace $X\subset H$ and pick $h\in X$, $h\neq 0$. It suffices to prove that any function in $H$ can be obtained as a finite linear combination of $G$-translates of $h$.
We have seen before that every function in $H$ is a finite linear combination of $G$-translates of $f$, therefore it suffices to show that $f$ is a finite linear combination of $G$-translates of $h$. Up to translating, we may assume $h(o)\neq 0$. If we average $h$ over $K$ (with respect to the Haar measure), we get a radial function that has to be a nonzero multiple of $f$. The operation of averaging over $K$ is the same as applying a finite linear combination of $G$-translates (actually $K$-translates), since the function $h$ is left-invariant under $G(B_{n}(o))$ for some $n$ and the quotient $K/G(B_{n}(o))=G(o)/G(B_{n}(o))$ is finite.
@@ 5 Preserved sesquilinear forms
We first recall that, due to Proposition 3.3, super-cuspidal and special representations (which are unitary) cannot preserve any nondegenerate sesquilinear form of index 1$\geq$ 1≥ 1. This means that to prove Theorem 2.3, it suffices to consider the case of spherical representations.
Therefore, let $(\pi,H)$ be an irreducible spherical representation of $G$ that is not 111-dimensional. We apply Theorem 4.6 and find that there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that $H$ is the space of functions $h:V\rightarrow\mathbb{K}$ that are left-invariant under some open compact subgroup of $G$ and satisfy $Lh=\alpha h$, and $\pi$ is the natural left action of $G$ on this space.
We split the space $H$ into a direct sum of $K$-invariant subspaces. First, we define $H_{0}$ as the 111-dimensional space generated by the unique (up to scalar multiple) $K$-invariant $f\in H$. For $v\in V$, we define the set $N(v)$ as in the proof of Theorem 4.6: $N(v)$ is the set of neighbors of $v$ that are further away from the origin than $v$ (or the complete set of neighbors if $v=o$). We define the space $\widetilde{H}_{1}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{1}(o))$, and then $H_{1}$ such that $\widetilde{H}_{1}=H_{0}\oplus H_{1}$, in this way:
@@
It is clear that $H_{0}$ is $K$-irreducible, since it is 111-dimensional. We show that $H_{1}$ is also $K$-irreducible. Since $K$ acts transitively on $\partial T$, it also acts transitively on $S_{1}$. Moreover, if $v\in S_{1}$ the stabilizer of $v$ in $G$ acts transitively on $\partial T$. This implies that its intersection with $K=G(o)$ acts transitively on $S_{1}\setminus\{v\}$, which (since $v$ is generic) is the same as saying that the action of $K$ on $S_{1}$ is doubly transitive. Now, let $X$ be any nonzero $K$-invariant subspace of $H_{1}$ and let $h\in X$ be nonzero. By averaging $h$ over $K\cap G(v)$, where $v\in S_{1}$ satisfies $h(v)\neq 0$, we get a nonzero scalar multiple of a function with value 111 at $v$ and $-\frac{1}{d-1}$ at all other vertices in $S_{1}$. Since $K$-translates of this function span $H_{1}$, we must have $X=H_{1}$.
@@
For $v\in V$, $v\neq o$, we define the cone $C_{v}$ as the set of vertices $w$ such that the unique path between $o$ and $w$ passes through $v$. We define the anti-cone $D_{v}:=(V\setminus C_{v})\cup\{v\}$. As $G$ has the independence property, it follows from Proposition 11 of [1] (applied to the subtree $B_{n-1}(o)$) that
@@
Clearly, all factors commute. If $X$ is a $K$-irreducible subspace of $H_{n}$ and $h\in X$ is nonzero, we can choose $v\in S_{n-1}$ such that $h(w)\neq 0$ for at least one $w\in N(v)$. We average $h$ across $\prod_{z\in S_{n-1}\setminus\{v\}}G(D_{v})$ and get a function that is unchanged on $N(v)$ and zero on $N(z)$ for every $z\in S_{n-1}\setminus\{v\}$. By the transitivity of $K$ on $S_{n-1}$ we can then get such a function for any $v\in S_{n-1}$. In this way, we get $|S_{n-1}|$ linearly independent functions in $X$, which proves our claim.
@@
To continue the proof of Theorem 2.3, we fix a $G$-invariant sesquilinear form $B$ on $H$. It follows from the Schur Lemma and Corollary 5.2 that $B$ must be the sum (for $n\in\mathbb{N}$) of $K$-invariant sesquilinear forms $B_{n}$ on $H_{n}$. This implies that if $\mathcal{A}$ and $\mathcal{B}$ are disjoint subsets of $\mathbb{N}$ then every function in $\bigoplus_{n\in\mathcal{A}}{H_{n}}$ has to be orthogonal to every function in $\bigoplus_{n\in\mathcal{B}}{H_{n}}$. We can now show the following:
@@
If the representation preserves the finite-index sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
@@
If we consider $g\in G$ that sends $o$ to $v$ and $v$ to $o$, we have $\pi(g)f=\alpha f+(1-\alpha^{2})h$ and $\pi(g)h=f-\alpha h$. The latter implies
$B(f,f)=\left(1-|\alpha|^{2}\right)B(h,h).$
Observe that $B(f,f)\neq 0$, because otherwise the form would be degenerate ($f$ is orthogonal to all functions in $\bigoplus_{n\geq 1}{H_{n}}$). Hence the last relation implies $B(h,h)\neq 0$.
As $B(f,h)=0$ it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(\alpha f+(1-\alpha^{2})h,f-\alpha h)=0$, and therefore
$\overline{\alpha}B(f,f)-(1-\overline{\alpha}^{2})\alpha B(h,h)=0.$

### v2

@@ 4 An explicit form for spherical representations
Proposition 4.1.
For any irreducible representation $(\pi,H)$ of $G$ that preserves a sesquilinear form of finite index $q$, the space $H^{K}$ of $K$-fixed vectors is at most $1$-dimensional.
@@
To prove our proposition, we follow the proof of Proposition 5.4 in [undefu]. As in that proof, we observe that this algebra is commutative and symmetric, the latter meaning that the adjoint of every element of $\mathcal{A}$ (with respect to the sesquilinear form preserved by $\pi$) is again in $\mathcal{A}$. Since in our case $g^{-1}\in KgK$ for every $g\in G$, a stronger statement holds: every element of $\mathcal{A}$ is self-adjoint.
@@
As in [undefu] we see that $\mathcal{A}$ acts irreducibly on $H^{K}$ (which we assume to be non-trivial), we see that the restriction of the sesquilinear form on $H$ to $H^{K}$ is non-degenerate (say of index $q^{\prime}\geq 0$), and we use the fact, due to Naimark ([undefw], Corollary 2) that a commutative symmetric algebra of operators of a complex Hilbert space endowed with a nondegenerate sesquilinear form of finite index $q^{\prime}\geq 1$ preserves a non-negative subspace $L$ of finite dimension. This is also true in the case $q^{\prime}=0$, due to the Schur Lemma, as explained in the addendum to [undefu].
@@
In the case $\mathbb{K}=\mathbb{C}$, the result in [undefw] applied to the representation of $\mathcal{A}$ on $H^{K}$ ensures that $\mathcal{A}$ preserves a finite-dimensional subspace $L$ of $H^{K}$, and as the representation is irreducible it must be $L=H^{K}$. Since $\mathcal{A}$ is commutative, every element of $\mathcal{A}$ is an endomorphism of the representation, so the Schur Lemma implies that $\mathcal{A}\cong\mathbb{C}$ and therefore $H^{K}$ is $1$-dimensional.
@@
In the case $\mathbb{K}=\mathbb{R}$, we apply the result in [undefw] to the complexification of the representation of $\mathcal{A}$ on $H^{K}$ (defined as $H^{K}\otimes_{\mathbb{R}}\mathbb{C}$). We find that there exists a finite-dimensional complex vector space $L$ preserved under $\mathcal{A}\otimes\mathbb{C}$ and thus under $\mathcal{A}$. This implies that there exists a finite-dimensional real vector space $L^{\prime}\subseteq H^{K}$ preserved under $\mathcal{A}$, which must coincide with $H^{K}$ by irreducibility. The representation preserves a positive definite bilinear form on $L^{\prime}=H^{K}$, since the restriction of the sesquilinear form to $L^{\prime}$ is non-negative and non-degenerate. If we apply the Schur Lemma, we see that $\mathcal{A}$ is a finite-dimensional commutative division algebra over $\mathbb{R}$ and therefore it is isomorphic to $\mathbb{R}$ or $\mathbb{C}$. In the latter case, there exists $A\in\mathcal{A}$ such that $A^{2}=-Id$, which implies, since each $A\in\mathcal{A}$ is self-adjoint, that for every $v\in H^{K}$ it holds
@@
contradicting the fact that the form is positive definite. Therefore, it must be $\mathcal{A}\cong\mathbb{R}$, which implies that $H^{K}$ is $1$-dimensional.
In the case $\mathbb{K}=\mathbb{H}$, we apply the result in [undefw] to the complex representation of $\mathcal{A}$ on $H^{K}$ seen as a complex vector space and find that it preserves a finite-dimensional complex subspace $L$. The quaternionic vector space that it generates is finite-dimensional, it must coincide with $H^{K}$ by irreducibility, and the sesquilinear form on it is positive definite. Applying the Schur Lemma, we again see that $\mathcal{A}\cong\mathbb{R}$ or $\mathcal{A}\cong\mathbb{C}$. The case $\mathcal{A}\cong\mathbb{C}$ can be excluded with the same argument as before, and we can conclude that $H^{K}$ is $1$-dimensional.
Proposition 4.1 implies that there exists a nonzero function $f\in H$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$.
@@
The function $f$ is an eigenfunction of the Laplacian, i. e. there exists some $\alpha\in\mathbb{K}$ such that $Lf=f\cdot\alpha$.
@@
Since the space of radial functions in $H$ (which is the same as the space of $K$-invariant functions in $H$) is $1$-dimensional and both $f$ and $Af$ are in $H$, we must have $Af=f\cdot\alpha$, and hence $Lf=f\cdot\alpha$, for some $\alpha\in\mathbb{K}$ (recall that we intend the scalar multiplication to be on the right if $\mathbb{K}=\mathbb{H}$).
It is immediate to see, from Observation 4.4, that the radial function $f$ is uniquely determined by $f(0)$ and by the property $Lf=f\cdot\alpha$. In particular, if $f(0)=0$ then $f$ is identically zero; therefore, we might assume $f(0)\neq 0$, and up to multiplying by a scalar $f(0)=1$. We can then calculate $f(1)=\alpha$ and $f(2)=\frac{1}{d-1}(df(1)\cdot\alpha-f(0))=\frac{(\alpha^{2}d)-1}{d-1}$. In general, $f(n)$ is a polynomial in $\alpha$ with real coefficients for every $n$. This implies that even in the case $\mathbb{K}=\mathbb{H}$, $\alpha$ commutes with $f(n)$ for every $n$, therefore $f$ satisfies $Lf=\alpha\cdot f$, where the notation indicates that each value of $f$ is multiplied by $\alpha$ on the left. This is particularly convenient since the space $\{h\in C:Lh=\alpha\cdot h\}$ is a right $\mathbb{K}$-vector subspace of $C$.
If $\alpha=1$, the function $f$ is constant on $V$ and thus $G$-invariant. Hence $H$ is generated by $f$ and the representation $(\pi,H)$ is the trivial representation. If $\alpha=-1$, the function $f$ has the same value $f(0)$ at every vertex at even distance from $o$ and the value f(0)-f(0) at every vertex at odd distance from $o$. An element of $G$ sends $f$ to itself if it moves vertices by an even distance, and to f-f otherwise. The space $H$ is again $1$-dimensional, and the representation $(\pi,H)$ is the $1$-dimensional character of order $2$; its kernel is the subgroup that moves vertices by an even distance.
We are left with the general case $\alpha\neq\pm 1$. This implies $f(2)\neq f(0)$: otherwise, we would have $(\alpha^{2}d)-1=d-1$ and thus $\alpha=\pm 1$.
@@
For any irreducible spherical representation $(\pi,H)$ of $G$ that is not $1$-dimensional, there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that the representation is equivalent to the left action of $G$ on the subspace of $C$ defined by $\{h\in C:Lh=\alpha\cdot h\}$. Conversely, for every $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, there is an irreducible spherical representation of $G$ that results from the construction.
@@
We start by picking a general irreducible spherical representation of $G$ and using the above results. We know that there exist $\alpha\in\mathbb{K}$ and a radial $f\in H\subset C$ with $f(0)=1$ such that $Lf=\alpha\cdot f$ and the subspace of $K$-invariant functions in $H$ is the space of scalar multiples of $f$. We can assume $\alpha\neq\pm 1$, and thus $f(2)\neq f(0)$: otherwise, as we have seen, the representation is $1$-dimensional.
Since the representation is irreducible, $H$ is generated (as a vector space) by $\{\pi(g)f:g\in G\}$. A function $\pi(g)f$ is radial with respect to $go$ and satisfies $L\pi(g)f=\alpha\cdot\pi(g)f$ (it is easily seen that both the Laplacian and the left multiplication by $\alpha$ commute with the action of $G$). As the property $Lf=\alpha\cdot f$ is preserved under linear combinations, we have $H\subseteq\{h\in C:Lh=\alpha\cdot h\}$.
To obtain the other inclusion, we pick a generic $h\in C$ such that $Lh=\alpha\cdot h$ and show that it is in $H$. Since $H$ is $G$-invariant and $f\in H$, it is enough to show that we can express $h$ as a finite linear combination of translates (by elements of $G$) of $f$.
For every $v\in V$, let $N(v)$ be the set of neighbors of $v$ that are further away from the origin than $v$ (if $v=o$ we take the full set of neighbors). To build $h$ as a finite linear combination of translates of $f$, we start with the function $F=f\cdot h(o)$, which agrees with $h$ at the origin (recall that $f(0)=1$). As $LF(o)=\alpha\cdot F(o)=\alpha\cdot h(o)=Lh(o)$, the sum of $F$ over $N(o)$ agrees with $h$. To make the two functions agree at every vertex in $N(o)$, we add to $F$ the function
$\sum_{w\in N(o)}{\pi(g_{w})f\cdot(f(0)-f(2))^{-1}(h(w)-F(w))},$
where $g_{w}\in G$ sends $o$ to $w$. The sum of the coefficients is zero and this guarantees that $F$ is unchanged at $o$. At every $w\in N(o)$ the function changes by
$f(0)(f(0)-f(2))^{-1}(h(w)-F(w))$
@@
$\sum_{z\in N(o)\setminus\{w\}}{f(2)(f(0)-f(2))^{-1}(h(z)-F(z))}=-f(2)(f(0)-f(2))^{-1}(h(w)-F(w))$
due to the contribution from $\pi(g_{z})f$ for $z\neq w$. This adds up to h(w)F(w)h(w)-F(w), so we have a new function $F$ that agrees with $H$ on the ball $B_{1}(o)$.
We can move forward by applying the same procedure to each of the sets $N(w)$ with $w\in N(o)$: this will allow to update the function $F$ so that it agrees with $h$ inside $N(w)$ without changing the values inside the ball $B_{1}(o)$, and ultimately get a function that agrees with $h$ on the ball $B_{2}(o)$. We can continue inductively with $B_{3}(o)$ and so on. After finitely many steps $F$ and $h$ will agree everywhere: $h$ is left-invariant under some open compact subgroup that contains $G(B_{n}(o))$ for some $n$ (Proposition 2.2(a)), so it is uniquely determined by the property $Lh=\alpha\cdot h$ and the values inside $B_{n}(o)$, which means that it suffices to have $F$ and $h$ agree on $B_{n}(o)$.
We have shown that the space $H$ coincides with $\{h\in C:Lh=\alpha\cdot h\}$, so the representation from which we started is equivalent to the left action of $G$ on that space.
For the converse, we wish to prove that if we start from a generic $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, and consider the space $\{h\in C:Lh=\alpha\cdot h\}$ with the left action of $G$, we have a spherical irreducible representation. We know that there is a radial (hence $K$-invariant) nonzero function in this space, so we are only left to prove irreducibility. Suppose there is a nonzero $G$-invariant subspace $X\subset H$ and pick $h\in X$, $h\neq 0$. It suffices to prove that any function in $H$ can be obtained as a finite linear combination of $G$-translates of $h$.
We have seen before that every function in $H$ is a finite linear combination of $G$-translates of $f$, therefore it suffices to show that $f$ is a finite linear combination of $G$-translates of $h$. Up to translating, we may assume $h(o)\neq 0$. If we average $h$ over $K$ (with respect to the Haar measure), we get a radial function that has to be a nonzero scalar multiple of $f$. The operation of averaging over $K$ is the same as applying a finite linear combination of $G$-translates (actually $K$-translates), since the function $h$ is left-invariant under $G(B_{n}(o))$ for some $n$ and the quotient $K/G(B_{n}(o))=G(o)/G(B_{n}(o))$ is finite.
@@ 5 Preserved sesquilinear forms
We first recall that, due to Proposition 3.3, super-cuspidal and special representations (which are unitary) cannot preserve any nondegenerate sesquilinear form of index $\geq 1$. This means that to prove Theorem 1.2, it suffices to consider the case of spherical representations.
Therefore, let $(\pi,H)$ be an irreducible spherical representation of $G$ that is not $1$-dimensional. We apply Theorem 4.6 and find that there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that $H$ is the space of functions $h:V\rightarrow\mathbb{K}$ that are left-invariant under some open compact subgroup of $G$ and satisfy $Lh=\alpha\cdot h$, and $\pi$ is the natural left action of $G$ on this space.
We split the space $H$ into a direct sum of $K$-invariant subspaces. First, we define $H_{0}$ as the $1$-dimensional space generated by the unique (up to scalar multiple) $K$-invariant $f\in H$. For $v\in V$, we define the set $N(v)$ as in the proof of Theorem 4.6: $N(v)$ is the set of neighbors of $v$ that are further away from the origin than $v$ (or the complete set of neighbors if $v=o$). We define the space $\widetilde{H}_{1}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{1}(o))$, and then $H_{1}$ such that $\widetilde{H}_{1}=H_{0}\oplus H_{1}$, in this way:
@@
It is clear that $H_{0}$ is $K$-irreducible, since it is $1$-dimensional. We show that $H_{1}$ is also $K$-irreducible. Since $K$ acts transitively on $\partial T$, it also acts transitively on $S_{1}$. Moreover, if $v\in S_{1}$ the stabilizer of $v$ in $G$ acts transitively on $\partial T$. This implies that its intersection with $K=G(o)$ acts transitively on $S_{1}\setminus\{v\}$, which (since $v$ is generic) is the same as saying that the action of $K$ on $S_{1}$ is doubly transitive. Now, let $X$ be any nonzero $K$-invariant subspace of $H_{1}$ and let $h\in X$ be nonzero. By averaging $h$ over $K\cap G(v)$, where $v\in S_{1}$ satisfies $h(v)\neq 0$, we get a nonzero scalar multiple of a function with value $1$ at $v$ and $\frac{1}{d-1}$ at all other vertices in $S_{1}$. Since $K$-translates of this function span $H_{1}$, we must have $X=H_{1}$.
@@
For $v\in V$, $v\neq o$, we define the cone $C_{v}$ as the set of vertices $w$ such that the unique path between $o$ and $w$ passes through $v$. We define the anti-cone $D_{v}:=(V\setminus C_{v})\cup\{v\}$. As $G$ has the independence property, it follows from Proposition 11 of [undef] (applied to the subtree $B_{n-1}(o)$) that
@@
Clearly, all factors commute. If $X$ is a $K$-irreducible subspace of $H_{n}$ and $h\in X$ is nonzero, we can choose $v\in S_{n-1}$ such that $h(w)\neq 0$ for at least one $w\in N(v)$. We average $h$ across $\prod_{z\in S_{n-1}\setminus\{v\}}G(D_{z})$ and get a function that is unchanged on $N(v)$ and zero on $N(z)$ for every $z\in S_{n-1}\setminus\{v\}$. By the transitivity of $K$ on $S_{n-1}$ we can then get such a function for any $v\in S_{n-1}$. In this way, we get $|S_{n-1}|$ linearly independent functions in $X$, which proves our claim.
@@
To continue the proof of Theorem 1.2, we fix a $G$-invariant nondegenerate sesquilinear form $B$ on $H$. It follows from the Schur Lemma and Corollary 5.2 that $B$ must be the sum (for $n\in\mathbb{N}$) of $K$-invariant sesquilinear forms $B_{n}$ on $H_{n}$. This implies that if $\mathcal{A}$ and $\mathcal{B}$ are disjoint subsets of $\mathbb{N}$ then every function in $\bigoplus_{n\in\mathcal{A}}{H_{n}}$ must be orthogonal to every function in $\bigoplus_{n\in\mathcal{B}}{H_{n}}$. We can now show the following:
@@
If the representation preserves the nondegenerate sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
@@
If we consider $g\in G$ that sends $o$ to $v$ and $v$ to $o$, we have $\pi(g)f=f\cdot\alpha+h\cdot(1-\alpha^{2})$ and $\pi(g)h=f-h\cdot\alpha$. The latter implies
$B(f,f)=B(h,h)\left(1-|\alpha|^{2}\right).$
Observe that $B(f,f)\neq 0$, because otherwise the form would be degenerate ($f$ is orthogonal to all functions in $\bigoplus_{n\geq 1}{H_{n}}$). Hence, the last relation implies $B(h,h)\neq 0$.
As $B(f,h)=0$ it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(f\cdot\alpha+h\cdot(1-\alpha^{2}),f-h\cdot\alpha)=0$, and therefore
$B(f,f)\overline{\alpha}-B(h,h)(1-\overline{\alpha}^{2})\alpha=0.$